# Classical and quantum beam dynamics simulation of the RF photoinjector test bench

### v3

@@ 5 Conclusions
We have performed classical and quantum beam-dynamics simulations of the S-band RF photoinjector test bench now being commissioned at JINR. Using realistic electromagnetic field maps and ASTRA tracking, we have showed that the injector can operate in a stable ultralow-charge regime ($Q\approx 0.63$ pC), where space-charge forces are weak and the emittance evolution is dominated by RF-induced correlations. Under these conditions, a Carlsten-type emittance-compensation scheme was implemented and optimized. Although the solenoid does not reduce the emittance below its intrinsic value, it suppresses RF-driven phase-space distortions and provides a reproducible operating point with a final normalized emittance of $\approx 3.2\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$.
The low-charge operating regime has enabled a quantum-mechanical analysis of single-electron states. We have studied the evolution of Laguerre–Gaussian wave packets with the realistic OAM values under three propagation scenarios: free drift, acceleration in a uniform DC field, and acceleration in the standing-wave RF field of the S-band gun. In both accelerating cases, the rapid growth of electron momentum has substantially suppressed transverse spreading compared to free-space propagation, while preserving the characteristic radial structure of the OAM-carrying states.
Overall, the simulations show that the present injector configuration satisfies the beam-quality requirements for experiments on the photoemission-driven generation and acceleration of vortex electrons in the MeV energy range. The photoinjector has been assembled, and the first photoemitted electron beam has been produced. Commissioning of the RF gun is ongoing, with work underway to reach the design input power of 6 MW. In parallel, a phase-locking system for synchronizing the UV laser pulses with the RF field is being developed. All preparations for experiments on photoemission-based generation of relativistic vortex electrons have been completed: diffractive optical elements for producing ultraviolet twisted light beams with orbital angular momentum up to $\ell=64\hbar$ have been fabricated and tested. Future studies will extend the present modelling by including collective interactions and decoherence mechanisms to obtain a complete description of structured electron beams under realistic experimental conditions. The check of the electron beams vorticity can be made by diffraction on a triangle apperture, as has been discussed in Ref. [Maksimov25].
@@ Acknowledgments
The authors express their gratitude to A.V. Vukolov, M.V. Shevelev, and G.K. Sizykh for their significant assistance in developing the beam emittance diagnostics. We also thank E.A. Yakushev, V.V. Glagolev, D. V. Naumov, A.S. Zhemchugov, and the management of the Laboratory of Nuclear Problems for supporting the project on behalf of JINR. Special thanks are due to K.V. Cherepanov and N.E. Sheremet for their help with quantum dynamics calculations, and to S.S. Baturin for valuable consultations.
This study was supported by the Russian Science Foundation (Project No. 23-62-10026) [RSF].

### v4

@@ 5 Conclusions
We have performed classical and quantum beam-dynamics simulations of the S-band RF photoinjector test bench now being commissioned at JINR. Using realistic electromagnetic field maps and ASTRA tracking, we have showed that the injector can operate in a stable ultralow-charge regime ($Q\approx 0.63$ pC), where space-charge forces are weak and the emittance evolution is dominated by RF-induced correlations. In this regime, the compensation solenoid serves mainly to control the transverse envelope and to regulate the magnetization introduced at the cathode. While the solenoid does not reduce the emittance below its intrinsic value, it suppresses RF-driven phase-space distortions and provides a reproducible operating point with a final normalized emittance of $\approx 2.08\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$.
The low-charge operating regime enables the investigation of single-electron states and their quantum properties. We have studied the evolution of Laguerre–Gaussian wave packets with the realistic OAM values under three propagation scenarios: free drift, acceleration in a uniform DC field, and acceleration in the standing-wave RF field of the S-band gun. In both accelerating cases, the rapid growth of electron momentum has substantially suppressed transverse spreading compared to free-space propagation, while preserving the characteristic radial structure of the OAM-carrying states.
Overall, the simulations show that the present injector configuration satisfies the beam-quality requirements for experiments on the photoemission-driven generation and acceleration of vortex electrons in the MeV energy range. The photoinjector has reached first beam, and commissioning toward the design RF input power of 6 MW is ongoing. In combination with a dedicated RF–laser synchronization system and tested diffractive optics for generating ultraviolet vortex beams up to $\ell=64\hbar$, all prerequisites for photoemission-based generation of relativistic vortex electrons are in place. Future studies will extend the present modelling by including collective interactions and decoherence mechanisms to obtain a complete description of structured electron beams under realistic experimental conditions. The check of the electron beams vorticity can be made by diffraction on a triangle apperture, as has been discussed in Ref. [29].
@@ Acknowledgments
The authors express their gratitude to A.V. Vukolov, M.V. Shevelev, and G.K. Sizykh for their significant assistance in developing the beam emittance diagnostics. We also thank E.A. Yakushev, V.V. Glagolev, D. V. Naumov, A.S. Zhemchugov, and the management of the Laboratory of Nuclear Problems for supporting the project on behalf of JINR. Special thanks are due to K.V. Cherepanov and N.E. Sheremet for their help with quantum dynamics calculations, and to S.S. Baturin for valuable consultations. The authors acknowledge A. V. Afanasyev for his valuable assistance in the assembly of the experimental setup.
This study was supported by the Russian Science Foundation (Project No. 23-62-10026) [37].
Appendix A RF phase scan
Figure 14: Simulated dependence of the final normalized transverse emittance on the injection RF phase $\varphi_{0}$. The operational working region $\varphi_{0}\approx 179^{\circ}$$181^{\circ}$.
Appendix B Beam dynamics without solenoid
Appendix C Particle tracking